# Europa: A Legal Multilingual Keyphrase Generation Dataset

### v1

@@ 1 Introduction
The scope of this work is to assess to what extent keyphrases can be automatically generated in the legal domain. Our contributions are the following : (1) We collected and curated Europa, the first open benchmark for legal KPG (keyphrase generation), spanning 24 languages and extracted from real-world European judgments; (2) We provide in-depth analysis of our corpus, highlighting its differences compared to other existing corpora; (3) We report performances achieved by multilingual generative models on this benchmark and point out areas where performances could be improved.  [^1] Upon acceptance of our work, the URL to the public GitHub repository with the complete dataset, scripts, and models will be provided in the camera-ready version of the paper.
@@ 2.1 Keyphrase Extraction and Generation
While research on KPG initially used seq-to-seq models in the form of RNNs Meng et al. (2017, 2019, 2021); Yuan et al. (2020a), modern KPG uses mostly fine-tuning of pre-trained transformer models (Vaswani et al., 2017), such as BART (Lewis et al., 2020), which showed itself to be as efficient as previous, more complex RNN models Chowdhury et al. (2022). Current English state-of-the-art results are currently obtained by using BART models which are further pre-trained for scientific KPG (KeyBART by Kulkarni et al. (2022)) and SciBART by Wu et al. (2022)). KPG has been conducted on limited domains, such as academic and STEM papers (Hulth, 2003a; Nguyen and Kan, 2007; Kim et al., 2010; Meng et al., 2017; Krapivin et al., 2009) or news Gallina et al. (2019); Koto et al. (2022), although the need for KPG extends beyond those domains.
@@ 2.2 Keyphrase Generation in the Context of Legal Domain
Input text (fr)
La demande de décision préjudicielle porte sur l’interprétation de l’article 48 du règlement (CEE) n° 1408/71 du Conseil, du 14 juin 1971, relatif à l’application des régimes de sécurité sociale aux travailleurs salariés, aux travailleurs non salariés et aux membres de leur famille qui se déplacent à l’intérieur de la Communauté (JO L 149, p. 2) […]
KPs (fr)
Assurance vieillesse – Travailleur ressortissant d’un État membre – Cotisations sociales – Périodes différentes – États membres différents – Calcul des périodes d’assurance – Demande de pension – Résidence dans un État tiers
Input text (de)
Das Vorabentscheidungsersuchen betrifft die Auslegung von Art. 48 der Verordnung (EWG) Nr. 1408/71 des Rates vom 14. Juni 1971 zur Anwendung der Systeme der sozialen Sicherheit auf Arbeitnehmer und Selbständige sowie deren Familienangehörige, die innerhalb der Gemeinschaft zu- und abwandern (ABl. L 149, S. 2) […]
KPs (de)
Altersversicherung – Arbeitnehmer mit der Staatsangehörigkeit eines Mitgliedstaats – Sozialbeiträge – Unterschiedliche Zeiten – Unterschiedliche Mitgliedstaaten – Berechnung der Versicherungszeiten – Rentenantrag – Wohnort in einem Drittland
Input text (en)
This reference for a preliminary ruling concerns the interpretation of Article 48 of Council Regulation (EEC) No 1408/71 of 14 June 1971 on the application of social security schemes to employed persons, to self-employed persons and to members of their families moving within the Community (OJ, English Special Edition 1971 (II), p. 416) […]
KPs (en)
Old-age insurance – Worker who is a national of a Member State – Social security contributions – Separate periods – Different Member States – Calculation of periods of insurance – Application for a pension – Residence in a non-Member State
@@ 3.2 Chronological Split
It is a common practice to randomly split data in a NLP task Gorman and Bedrick (2019). However, training and evaluating a model on data from overlapping time periods with similar distributions fails to assess the actual model’s temporal generalization (Lazaridou et al., 2021). This is why legal NLP generally uses a chronological split of documents instead of combining random shuffle with random split (Chalkidis et al., 2019; Medvedeva et al., 2021; Chalkidis et al., 2021). In our case, we tried both splits with a mBART50 model. On the test set, the performance in terms of $F1@M$ for present keyphrases differs by 14.1 percentage points (11.1 for absent keyphrases) in favour of random split. Results in Table 2 confirm that random split, by ignoring real-world temporal concept drifts, tends to overestimate true performance. This is consistent with Søgaard et al. (2021); Mu et al. (2023). Therefore, we choose a chronological split for a proper assessment: The training set covers judgments from 1957 to 2010 (131 076 instances), the validation those from 2011 to 2015 (63 373 instances), and the test set the ones from 2016 to 2023 (90 508 instances). We will make available a random split version of the dataset for those who wish it. Full details about documents distribution across these splits are shown in Appendix F.
@@ 4 Dataset Analysis
As shown in Table 3, compared to other previous works and KPG benchmarks, Europa covers more languages and is highly positioned in matters of number of instances, ratio of absent keyphrases, and average input length.
@@ 5.1 Evaluation Protocol
1k
8k
1k
8k
French
33.1
34.4
8.7
9.6
German
33.5
35.8
4.9
5.0
English
29.5
31.7
5.0
5.1
Italian
30.5
33.3 3.7
3.6
Dutch
31.9
34.4
3.5
3.7
Greek
14.8
15.7
10.6
11.1
Danish
30.4
32.6 3.2
2.6
Portuguese
27.6
30.1
5.5
6.5
Spanish
29.2
31.8
4.8
4.8
Swedish
33.1
33.9
4.2
5.0
Finnish
27.5
28.9
11.6
13.1
Lithuanian
34.9
32.9
4.2
4.7
Estonian
29.5
31.0 4.7
4.5
Czech
26.4
29.5
13.7
14.6
Hungarian
14.3
15.3
10.4
11.3
Latvian
34.7
33.1
3.5
3.6
Slovene
25.6
29.2
11.1
11.8
Polish
26.3
29.7
11.7
13.2
Maltese
24.7
27.0 7.7
7.1
Slovak
25.2
24.9
15.7
16.0
Romanian
30.7
33.8
5.4
5.7
Bulgarian
29.0
29.2
3.4
3.5
Croatian
8.6
15.5 4.1
3.8
Irish
0.0
2.2
0.0
0.0
@@ 6 Results
Unsurprisignly, extractive model YAKE performs poorly and finds none of the absent keyphrases. With a fixed input length of 1024 tokens, the three mT5 variants dramatically underperform mBART50. This is surprising as mT5 covers twice as many languages than mBART50. Also, as mT5-large has 59% more parameters than mBART50, it would have been expected to outperform the latter, but the results reveal otherwise. When comparing mBART50 with mBART50-8k, the increase in the maximum context length brings significant improvement across all metrics. While an average gain of around 2% (for $F1@M$ over present phrases) may not seem significant, KPG evaluation often greatly underestimates the true performance of the models, due to the difficulty of correctly evaluating whether a keyphrase is correct or not (Wu et al., 2023). As such, the improvement shown by the mBART50-8k model is significant and reflected in the generated keyphrases, thus emphasizing the benefits in enlarging the maximum input length. Still, Table 5 shows that input context enlargement do not have uniform effects across languages. For instance, for present keyphrases, some languages get small gains (Greek, Swedish), and some degrade (Lithuanian, Latvian). For high-resource languages such as English and Italian, performance improves on present phrases, but seems stagnant on absent ones. For low-resource languages, Croatian has the most dramatic improvement on present phrases while Irish gets one score above 00 with mBART50-8k. However, the performance for these languages still lags behind that of moderate-resource ones. This is understandable as these languages have almost no training instances in our temporal split setting, thus revealing the difficulty of conducting KPG for unseen languages [^4] With a random split dataset, $F1@M$ achieved by mBART50 for Croation/Irish reaches 46.9/20.5 for present keyphrases and 13.9/10.2 for absent ones.
@@ 7 Analysis and Discussion
The improvement from mBART50 to mBART50-8k is significant and emphasizes the importance of larger input length. Less that 14% of all instances fit into a 1024-tokens context window. Around 58% do when that length reaches 8k tokens. Building models that can efficiently generate keyphrases from larger documents is therefore crucial in order to achieve further progress (the length reaches 9160 tokens on average, and 17k at the 90th percentile, with mBART50 tokenizer) [^5] We tried a sliding-window-based model that was not conclusive, thus the need to find better ways to capture context..
@@
YAKE
0.335
0.155
mT5-small
0.399
0.268
mT5-base
0.422
0.283
mT5-large
0.436
0.270
mBART50
0.519 0.442
mBART50-8k
0.548 0.479
@@ 8 Conclusion
In this paper, we present Europa, a novel multilingual keyphrase generation dataset in the legal domain. We believe this dataset can help alleviate two current shortcomings of the keyphrase generation task : the lack of data in domains other than STEM; and the lack of multilingual non-English datasets. We furthermore provide an analysis of Europa with key statistics, thus giving insights on the particularities of our dataset. Finally, we run multiple models in various settings in order to give an initial point of comparison for future works on Europa. Our corpus also highlights the need to efficiently capture larger input context, and will be a suitable testbed for models designed to do so.
@@
Low Resource Languages: The choice of a chronological split, though enabling a realistic KPG performance assessment with quasi-equal amount of test instances for each language (except Irish), results in dramatic differences in available training data across languages. This is particularly true for the latest official EU languages, such as Croatian and Irish. One possible approach for mitigating this issue in future works would be to complement the training data with other legal documents released by other EU institutions. For the time being, a random split version of our dataset will be made available for those who would like an identical distribution of languages across training, validation and test sets.
Computing Cost Scalability: The architecture of most of the models used here is derived from experiments in which in the text input length rarely exceeds 1024. This is a major caveat when deploying such models on real-world data with significantly longer documents, such as the legal documents in our dataset which often exceed several thousand tokens. Moreover, dealing with documents whose language is not supported by the tokenizer mechanically increases the number of tokens. This is due to the fact that running and training transformer models with higher maximum input sequences is costly both in terms of time and GPU memory. The computing burden was already emphasized by Aumiller et al. (2022) of legal summarization for EU documents and Sakiyama et al. (2023) for legal KPG in Portuguese.
@@ Acknowledgements
We would like to thank the European Union for granting us access to EUR-Lex data (http://eur-lex.europa.eu, © European Union, 1998-2023).
@@ Appendix B Languages Support per Model
Model
mBART50
mT5
Total lang.
52
101
Unsupported EU lang.
bg, da, el, ga, hu, mt, sk
hr
@@ Appendix C Number of Parameters per Model
Model # param.
@@ Appendix D Training and Inference Times
Model Num. of training epochs
Total training time
Total inference time
mT5-small 10
11h
7h
mT5-base 10
28h
9h
mT5-large* 5
53h
16h
mBART50 10
25h
31h
mBART50-8k* 5
61h
34h
@@ Appendix H Random Split
Still, to better allow practitioners to use data as they desire, we will release a random split. The training, validation and test sets have respectively 10 003, 3 391 and 4 439 judgments. These figures are the same as in the chronologically split dataset. However, the number of instances per set in the random split setting differs since each set has the same distribution in terms of languages. Consequently, each set contains 159 306, 53 943 and 71 708 instances respectively. The results achieved by mBART50 are shown in Table 15. As expected, results are overall much higher since all training, validation and testing sets have the same distribution in terms of languages and vocabulary. We want to emphasize once again however that due to the fast changing nature of legal NLP, these performances are only valid for a short time after the data collection.

### v2

@@ 1 Introduction
The scope of this work is to assess to what extent keyphrases can be automatically generated in the legal domain. Our contributions are the following: (1) We collected and curated Europa, the first open benchmark for legal KPG (keyphrase generation), spanning 24 languages and extracted from real-world European judgments. [^1] The dataset can be downloaded from the Hugging Face hub at https://huggingface.co/datasets/NCube/europa while evaluation scripts and model outputs are available at https://github.com/rali-udem/europa. (2) We provide in-depth analysis of our corpus, highlighting its differences compared to other existing corpora. (3) We report performances achieved by multilingual generative models on this benchmark and point out areas where performances could be improved.
@@ 2.1 Keyphrase Extraction and Generation
While research on KPG initially used seq-to-seq models in the form of RNNs Meng et al. (2017, 2019, 2021); Yuan et al. (2020a), modern KPG uses mostly fine-tuning of pre-trained transformer models (Vaswani et al., 2017), such as BART (Lewis et al., 2020), which showed itself to be as efficient as previous, more complex RNN models Chowdhury et al. (2022). Current English state-of-the-art results are currently obtained by using BART models which are further pre-trained for scientific KPG (KeyBART by Kulkarni et al. (2022) and SciBART by Wu et al. (2022)). KPG has been conducted on limited domains, such as academic and STEM papers (Hulth, 2003a; Nguyen and Kan, 2007; Kim et al., 2010; Meng et al., 2017; Krapivin et al., 2009) or news Gallina et al. (2019); Koto et al. (2022), although the need for KPG extends beyond those domains.
@@ 2.2 Keyphrase Generation in the Context of Legal Domain
Input text (fr) La demande de décision préjudicielle porte sur l’interprétation de l’article 48 du règlement (CEE) n° 1408/71 du Conseil, du 14 juin 1971, relatif à l’application des régimes de sécurité sociale aux travailleurs salariés, aux travailleurs non salariés et aux membres de leur famille qui se déplacent à l’intérieur de la Communauté (JO L 149, p. 2) […]
KPs (fr) Assurance vieillesse – Travailleur ressortissant d’un État membre – Cotisations sociales – Périodes différentes – États membres différents – Calcul des périodes d’assurance – Demande de pension – Résidence dans un État tiers
Input text (de) Das Vorabentscheidungsersuchen betrifft die Auslegung von Art. 48 der Verordnung (EWG) Nr. 1408/71 des Rates vom 14. Juni 1971 zur Anwendung der Systeme der sozialen Sicherheit auf Arbeitnehmer und Selbständige sowie deren Familienangehörige, die innerhalb der Gemeinschaft zu- und abwandern (ABl. L 149, S. 2) […]
KPs (de) Altersversicherung – Arbeitnehmer mit der Staatsangehörigkeit eines Mitgliedstaats – Sozialbeiträge – Unterschiedliche Zeiten – Unterschiedliche Mitgliedstaaten – Berechnung der Versicherungszeiten – Rentenantrag – Wohnort in einem Drittland
Input text (en) This reference for a preliminary ruling concerns the interpretation of Article 48 of Council Regulation (EEC) No 1408/71 of 14 June 1971 on the application of social security schemes to employed persons, to self-employed persons and to members of their families moving within the Community (OJ, English Special Edition 1971 (II), p. 416) […]
KPs (en) Old-age insurance – Worker who is a national of a Member State – Social security contributions – Separate periods – Different Member States – Calculation of periods of insurance – Application for a pension – Residence in a non-Member State
@@ 3.2 Chronological Split
It is a common practice to randomly split data in a NLP task Gorman and Bedrick (2019). However, training and evaluating a model on data from overlapping time periods with similar distributions fails to assess the actual model’s temporal generalization (Lazaridou et al., 2021). This is why legal NLP generally uses a chronological split of documents instead of combining random shuffle with random split (Chalkidis et al., 2019; Medvedeva et al., 2021; Chalkidis et al., 2021). In our case, we tried both splits with a mBART50 model. On the test set, the performance in terms of $F1@M$ for present keyphrases differs by 14.1 percentage points (11.1 for absent keyphrases) in favour of random split. Results in Table 2 confirm that random split, by ignoring real-world temporal concept drifts, tends to overestimate true performance. This is consistent with Søgaard et al. (2021); Mu et al. (2023). Therefore, we choose a chronological split for a proper assessment: The training set covers judgments from 1957 to 2010 (131 076 instances), the validation those from 2011 to 2015 (63 373 instances), and the test set the ones from 2016 to 2023 (90 508 instances). [^4] The temporal split is available at https://huggingface.co/datasets/NCube/europa and a random split version for those who wish it can be found at https://huggingface.co/datasets/NCube/europa-random-split Full details about documents distribution across these splits are shown in Appendix F.
@@ 4 Dataset Analysis
As shown in Table 3, compared to previous works and KPG benchmarks, Europa covers more languages and is highly positioned in matters of number of instances, ratio of absent keyphrases, and average input length.
@@ 5.1 Evaluation Protocol
1k 8k 1k 8k
French 33.1 34.4 8.7 9.6
German 33.5 35.8 4.9 5.0
English 29.5 31.7 5.0 5.1
Italian 30.5 33.3 3.7 3.6
Dutch 31.9 34.4 3.5 3.7
Greek 14.8 15.7 10.6 11.1
Danish 30.4 32.6 3.2 2.6
Portuguese 27.6 30.1 5.5 6.5
Spanish 29.2 31.8 4.8 4.8
Swedish 33.1 33.9 4.2 5.0
Finnish 27.5 28.9 11.6 13.1
Lithuanian 34.9 32.9 4.2 4.7
Estonian 29.5 31.0 4.7 4.5
Czech 26.4 29.5 13.7 14.6
Hungarian 14.3 15.3 10.4 11.3
Latvian 34.7 33.1 3.5 3.6
Slovene 25.6 29.2 11.1 11.8
Polish 26.3 29.7 11.7 13.2
Maltese 24.7 27.0 7.7 7.1
Slovak 25.2 24.9 15.7 16.0
Romanian 30.7 33.8 5.4 5.7
Bulgarian 29.0 29.2 3.4 3.5
Croatian 8.6 15.5 4.1 3.8
Irish 0.0 2.2 0.0 0.0
@@ 6 Results
Unsurprisignly, the YAKE extractive model performs poorly, finding none of the absent keyphrases. With a fixed input length of 1024 tokens, the three mT5 variants dramatically underperform mBART50. This is surprising as mT5 covers twice as many languages than mBART50. Also, as mT5-large has 59% more parameters than mBART50, it would have been expected to outperform the latter, but the results reveal otherwise. When comparing mBART50 with mBART50-8k, the increase in the maximum context length brings significant improvement across all metrics. While an average gain of around 2% (for $F1@M$ over present phrases) may not seem significant, KPG evaluation often greatly underestimates the true performance of the models, due to the difficulty of correctly evaluating whether a keyphrase is correct or not (Wu et al., 2023). As such, the improvement shown by the mBART50-8k model is significant and reflected in the generated keyphrases, thus emphasizing the benefits in enlarging the maximum input length. Still, Table 5 shows that input context enlargement does not have uniform effects across languages. For instance, for present keyphrases, some languages get small gains (Greek, Swedish), and some degrade (Lithuanian, Latvian). For high-resource languages such as English and Italian, performance improves on present phrases, but seems stagnant on absent ones. For low-resource languages, Croatian has the most dramatic improvement on present phrases while Irish gets one score above 00 with mBART50-8k. However, the performance for these languages still lags behind that of moderate-resource ones. This is understandable as these languages have almost no training instances in our temporal split setting, thus revealing the difficulty of conducting KPG for unseen languages. [^5] With a random split dataset, $F1@M$ achieved by mBART50 for Croation/Irish reaches 46.9/20.5 for present keyphrases and 13.9/10.2 for absent ones
@@ 7 Analysis and Discussion
The improvement from mBART50 to mBART50-8k is significant and emphasizes the importance of larger input length. Less that 14% of all instances fit into a 1024-tokens context window. Around 58% do when that length reaches 8k tokens. Building models that can efficiently generate keyphrases from larger documents is therefore crucial in order to achieve further progress (the length reaches 9160 tokens on average, and 17k at the 90th percentile, with mBART50 tokenizer). [^6] We tried a sliding-window-based model that was not conclusive, thus the need to find better ways to capture context.
@@
YAKE 0.335 0.155
mT5-small 0.399 0.268
mT5-base 0.422 0.283
mT5-large 0.436 0.270
mBART50 0.519 0.442
mBART50-8k 0.548 0.479
@@ 8 Conclusion
In this paper, we present Europa, a novel and open multilingual keyphrase generation dataset in the legal domain. We believe this dataset can help alleviate two current shortcomings of the keyphrase generation task: the lack of data in domains other than STEM; and the lack of multilingual non-English datasets. Our dataset is available at https://huggingface.co/datasets/NCube/europa. We furthermore provide an analysis of Europa with key statistics, thus giving insights on the particularities of our dataset. Finally, we run multiple models in various settings in order to give an initial point of comparison for future works on Europa. Our corpus also highlights the need to efficiently capture larger input context, and will be a suitable testbed for models designed to do so.
@@
Low Resource Languages: The choice of a chronological split, though enabling a realistic KPG performance assessment with quasi-equal amount of test instances for each language (except Irish), results in dramatic differences in available training data across languages. This is particularly true for the latest official EU languages, such as Croatian and Irish. One possible approach for mitigating this issue in future works would be to complement the training data with other legal documents released by other EU institutions. A random split version of our dataset is available for those who would like an identical distribution of languages across training, validation and test sets (https://huggingface.co/datasets/NCube/europa-random-split).
Computing Cost Scalability: The architecture of most of the models used here is derived from experiments in which the text input length rarely exceeds 1024. This is a major caveat when deploying such models on real-world data with significantly longer documents, such as the legal documents in our dataset which often exceed several thousand tokens. Moreover, dealing with documents whose language is not supported by the tokenizer mechanically increases the number of tokens. This is due to the fact that running and training transformer models with higher maximum input sequences is costly both in terms of time and GPU memory. The computing burden was already emphasized by Aumiller et al. (2022) of legal summarization for EU documents and Sakiyama et al. (2023) for legal KPG in Portuguese.
@@ Acknowledgements
We would like to thank the European Union for granting us access to EUR-Lex data (http://eur-lex.europa.eu, © European Union, 1998-2023). We are also grateful to Benjamin Cérat and Di Wu for their advice on some coding implementations.
@@ Appendix B Languages Support per Model
Model mBART50 mT5
Total lang. 52 101
Unsupported EU lang. bg, da, el, ga, hu, mt, sk hr
@@ Appendix C Number of Parameters per Model
{tblr}
column2 = r, m, cell12 = c, hline1,7 = -0.08em, hline2 = -0.05em, Model & # param.
@@ Appendix D Training and Inference Times
{tblr}
columneven = c, column3 = c, hline1,7 = -0.08em, hline2 = -0.05em, Model & Num. of training epochs Total training time Total inference time
mT5-small 10 11h 7h
mT5-base 10 28h 9h
mT5-large* 5 53h 16h
mBART50 10 25h 31h
mBART50-8k* 5 61h 34h
@@ Appendix H Random Split
Still, to better allow practitioners to use data as they desire, we released a random split (https://huggingface.co/datasets/NCube/europa-random-split). The training, validation and test sets have respectively 10 003, 3 391 and 4 439 judgments. These figures are the same as in the chronologically split dataset. However, the number of instances per set in the random split setting differs since each set has the same distribution in terms of languages. Consequently, each set contains 159 306, 53 943 and 71 708 instances respectively. The results achieved by mBART50 are shown in Table 15. As expected, results are overall much higher since all training, validation and testing sets have the same distribution in terms of languages and vocabulary. We want to emphasize once again however that due to the fast changing nature of legal NLP, these performances are only valid for a short time after the data collection.